Service Status

Grey Line

Follow us on twitter or Like our Facebook Page for live status updates.


IPv6 routing issues - 27/07/2012 08:45

The earlier IPv6 routing issues are now resolved.


IPv6 routing issues - 27/07/2012 07:30

We are aware of some intermittent IPv6 routing problems. Engineers are investigating.


Power Failure - 11/04/2012 18:49BST

The failed shared webserver has been restored to full service.


Power Failure - 11/04/2012 17:29BST

One of our shared webservers has not returned following the power outage, engineers are working on the problem.


Power Failure - 11/04/2012 16:09BST

All email and web services are restored to normal.


Power Failure - 11/04/2012 15:55BST

Telephone service is restored


Power Failure - 11/04/2012 15:39BST

Engineers switch UPS to bypass and Datacentre load taken by standby generators.


Power Failure - 11/04/2012 15:34BST

Lightning strike causes power & telephone failure


POP Server Emergency Maintenance - 04/09/2012 1500-1800BST

POP1 is undergoing emergency maintenance between 1500-1800BST. Email for this server may not be available during this window.


POP Server Update - 05/04/2012 1050

We had a brief outage this morning which has been now resolved.


POP Server Update - 05/04/2012 0610

The affected POP server was restored to full operation at 06:50 BST and all users should now have functioning E-mail. The backlog of E-mails sent overnight are currently being processed and will be delivered shortly. If you still have problems with E-mail, please contact us.


POP Server Emergency Maintenance - 04/04/2012 2130

A hardware failure occurred earlier on the pop1 POP/IMAP server which has affected all customers whose E-mail is handled by this machine. The hardware has been replaced and the machine is back in service only to allow data to be restored. This process is likely to take several hours during which time the server is not accepting new mails - however, all mails sent will be queued and will be delivered when the server is fully online again.


POP Server Emergency Maintenance - 04/04/2012 1915

Engineers are working to resolve a server issue


Reduced Functionality - 19/02/2012 0945-

All services have been restored. We are obviously investigating the length of time taken internally and with our third party suppliers.


Reduced Functionality - 19/02/2012 0810-

We are advised by our supplier that this should be resolved shortly.


Reduced Functionality - 18/02/2012 2100-

Some services are still affected, and we are still working with our supplier to resolve the issue.


Reduced Functionality - 18/02/2012 1645-

There are issues with rerouting around the problem; this has been escalated and we are in the process of relocating some affected services.


Reduced Functionality - 18/02/2012 1430-

We are still working with our supplier to resolve the ongoing issue, although this is taking longer than anticipated.


Reduced Functionality - 18/02/2012 1250-

Initial attempts to restore connectivity within our provider's network have been unsuccessful. Their engineers are working on an alternative strategy; we still expect a speedy resolution.


Reduced Functionality - 18/02/2012 1050-

We are continuing to work with our third party provider to resolve this issue. We hope to have a resolution soon.


Reduced Functionality - 18/02/2012 0930-

Due to failure in one of our third-party providers, part of our network is on a backup connection of lower capacity. Some customers may experiences slow or intermittent connection issues.

Our engineers are working with our third-party providers to bring this back up as soon as possible.

We apologise for any inconvenience caused.


Urgent Maintenance - 03/09/2011 1800 - 04/09/2011 0200

Affected customers: All hosting customers, broadband access customers and leased line access customers.

We are performing urgent maintenance in our Telehouse PoP to address an issue with the power distribution units in one of our main infrastructure racks. As a result, some items of equipment will briefly lose power as they are moved onto a temporary power supply whilst the faults are rectified, and then moved back onto the new supply.

We are also taking the opportunity of this power maintenance to replace some older network equipment that does not support IPv6, thus completing our planned IPv6 roll-out across the network.

Please do not call or raise a ticket about any problems encountered during this time unless you are still experiencing issues after we have sent out a completion E-mail.


14:00 - 22 Apr 2011 - Partial Infrastructure Outage London

A customer facing switch died in our telehouse hosting centre, engineers have now replaced and can confirm service has been restored.

Please accept our apologies for any inconvenience this may have caused.


15:00 - 11 Jan 2011 - ADSL Authentication Issue

Although we have not yet had an update from BT, the issue appears to be fixed and has been stable since 13:00hrs.

Once again please accept our apologies


02:00 hrs 11 Jan 2011 - ADSL Authentication Issue

We have been informed that there is a major outage at this time affecting several 21CN lines.
Our engineers are working along with BT to get this resolved on priority.
Unfortunately, we do not have any estimated time for resolution. However, we will keep this page updated.

Please accept our sincere apologies for the inconvenience caused.